The 11-N horizontal force acts on the handle of the socket wrench as shown. Determine the (a) moment of the 11-N force about Please note the changed values at the top!
Answer #1

ht r = 0.288m 160 140128 Cos 16+ 0.085k m. = 0.0771 +0.26917010864 8 = -11 cosibi+11 Gin 18° j 2 of = -10.574 i + 3.0328 N I